Problems 1-13: An appliance store recorded data for their customers who purchased warranties and had complaints about their appliances. The following data classifies the categories of complaints with whether the complaint occurred during or after the warranty. There were 100 complaints. Express your answers as decimals. If necessary, round to four decimal positions. Problems 14 and 15 are short answer, requiring justification.

                                   Electrical                       Mechanical                       Appearance

  

During Warranty             18                                  13                                      31

After Warranty                12                                  22                                       4

Fill in the blank by typing the correct answer.

6

If two complaints are selected randomly without replacement (first not replaced before second selected), find:

Probability that both happened during warranty =

7

If two complaints are selected randomly without replacement (first not replaced before second selected), find:

Probability that at least one happened after warranty =

8

If two complaints are selected randomly without replacement (first not replaced before second selected), find:

Probability that neither happened during warranty =

9

If two complaints are selected randomly without replacement (first not replaced before second selected), find:

Probability that the first was electrical and second was mechanical =

10

If two complaints are selected randomly with replacement (first is replaced before second selected), find:

Probability that both happened during warranty:  

11

If two complaints are selected randomly with replacement (first is replaced before second selected), find:

Probability that neither happened during warranty =

12

If a sample of four complaints is randomly selected without replacement, find:

Probability that all four happened after warranty =

13

If a sample of four complaints is randomly selected without replacement, find:

Probability that at least one happened during warranty =

Complete the short answer question in 500 characters using complete sentences:

14

Given that a claim happened during warranty, what is the most likely type of claim? Justify your answer in a sentence.

Type answer here.

Your answer should not have more than 500 characters.

15

Companies have to balance profit with reputation. Does it seem that this company has a potential issue with customer service even if it takes responsibility for “appearance” claims during warranty? What issues might a customer have with the appearance of an appliance after it is installed? Do you think a customer is likely to return an appliance for appearance reasons six months after it is in use? Explain in a short paragraph.
Answer #1
electrical mechancial appearance total
during 18 13 31 62
after 12 22 4 38
total 30 35 35 100

6 P(both happened during warranty) = 62/100 * 61/99 =                       0.3820  
                          
7) P(atleast one happened after warranty) = P(one happen during warranty)P(one happen after warranty)+P(both happen after warranty) = 62/100*38/99 + 38/100*62/99 + 38/100*31/99=                       0.5949  
                          
8) P(neither happen during warranty)=P(both happen after warranty) = 38/100*37/100=                       0.1406  
                          
9) P(first was electrical and second was mechanical )=30/100*35/99=                       0.1061  
                          
10) P(both happen during warranty) = 62/100*62/100=                       0.3844   [because problem is of replacement]
------------------------------

please revert if have doubts